# Training Language Models on Synthetic Edit Sequences Improves Code Synthesis

### v2

@@ Abstract
Software engineers mainly write code by editing existing programs. In contrast, large language models (LLMs) autoregressively synthesize programs in a single pass. One explanation for this is the scarcity of open-sourced edit data. While high-quality instruction data for code synthesis is already scarce, high-quality edit data is even scarcer. To fill this gap, we develop a synthetic data generation algorithm called LintSeq. This algorithm refactors existing code into a sequence of code edits by using a linter to procedurally sample across the error-free insertions that can be used to sequentially write programs. It outputs edit sequences as text strings consisting of consecutive program diffs. To test LintSeq, we use it to refactor a dataset of instruction + program pairs into instruction + program-diff-sequence tuples. Then, we instruction finetune a series of smaller LLMs ranging from 2.6B to 14B parameters on both the re-factored and original versions of this dataset, comparing zero-shot performance on code synthesis benchmarks. We show that during repeated sampling, edit sequence finetuned models produce more diverse programs than baselines. This results in better inference-time scaling for benchmark coverage as a function of samples, i.e. the fraction of problems “pass@k” solved by any attempt given “k” tries. For example, on HumanEval pass@50, small LLMs finetuned on synthetic edit sequences are competitive with GPT-4 and outperform models finetuned on the baseline dataset by +20% ($\pm$3%) in absolute score. Finally, we also pretrain our own tiny LMs for code understanding. We show that finetuning tiny models on synthetic code edits results in state-of-the-art code synthesis for the on-device model class. Our 150M parameter edit sequence LM matches or outperforms code models with twice as many parameters, both with and without repeated sampling, including Codex and AlphaCode.
@@ 1 Introduction
The successes of large language models (LLMs) are difficult to overstate. However, consistent and correct zero-shot generation in code synthesis remains out-of-reach for all but the largest models (Abdin et al., 2024; Groeneveld et al., 2024; Dubey et al., 2024). Compared to other reasoning tasks, this setting has two challenging properties, namely solutions are both structured and long-form.
@@
Figure 1: Code synthesis with LMs trained on synthetic code edit sequences. Left: An example generation from an LM trained to synthesize code as a stream of static-error-free edits. Right: Comparing zero-shot HumanEval coverage (in %) as a function of FLOPs for large external LLMs vs the smaller LMs that we instruction finetune in this paper. Repeatedly sampling from edit sequence LMs yields coding problem solutions that are competitive with GPT-4 and GPT-4-Omni, and have total cost similar to sampling once from the best open-source LLMs (see Appendix F.4).
@@
In contrast, LLMs are trained to autoregressively synthesize entire programs from scratch. This makes repeatedly editing a program with an LLM extremely expensive – current state-of-the-art, LLM-powered code editing tools like Cursor repeatedly prompt models to rewrite entire programs during every edit generation call (Sanger, 2024). LLM outputs also suffer from degrading quality as sequence lengths grow and exhibit limited diversity across samples (Chen et al., 2021; Li et al., 2022b; Roziere et al., 2023; Lozhkov et al., 2024). The consequence of these pathologies is that there does not exist a reliable trade-off between zero-shot generation quality and inference-time compute cost under the current paradigm of autoregressive code synthesis, particularly for smaller models.
@@
In this paper, we claim that these issues can be mitigated at the data-level by reparameterizing code synthesis as a sequential edit problem. Rather than training models for single-step generation of entire programs, we propose that models be trained to generate code by predicting code edit sequences.
@@
This objective has a major obstacle: while high-quality solution data for code synthesis is scarce, open-source edit data with good coverage over the distribution of all error-free code “diffs” is non-existent (Muennighoff et al., 2023). To address this, we introduce an algorithm titled “LintSeq” that re-factors existing programs into sequences of static error-free code edits. LMs trained on data generated with our algorithm output code edits that effect interdependent lines of a program. LintSeq is parameter-free. It consists of two phases: a backward sampling phase, which takes a source file as input and uses a static program verifier to sample sequences of error-free program states that begin with this file and end with empty programs; and a forward edit computation phase, which reverses each sequence of programs, employing the Unix diff (Thompson & Ritchie, 1975) operator to compute deltas between consecutive versions of the source file, and outputs edit sequences. The static program verifier used by LintSeq in its backward sampling phase is also referred to as a linter.
To test the impact of finetuning LMs to synthesize code with edit sequences via LintSeq instruction data, we conduct a series of experiments comparing these models to those finetuned on a standard version of the same instruction data. We evaluate LMs zero-shot on the code synthesis benchmarks HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) by computing coverage, the proportion of problems solved by any attempt, as a function of samples. We do not allow LMs to use chain-of-thought reasoning (Wei et al., 2022) during evaluation. Our results are five-fold:
@@
    Across models ranging in scale from 150M to 14B parameters, instruction finetuning LMs on edit sequences vs full programs improves the quality and diversity of synthesized code.
@@
    The improved diversity of samples means that pass@k performance increases smoothly as a function of inference-time compute, allowing for a better trade-off between the two.
@@
    Tiny edit sequence LMs have state-of-the-art performance in their model class (Table 1).
  4. 4.
    For smaller LLMs, repeatedly sampling from edit sequence models results in HumanEval coverage that is competitive with GPT-4 models and has similar cumulative cost to sampling once per problem from open-source LLMs like Llama 3.1 405B (Figures 1 and 4).
  5. 5.
    Ablating the linter from edit sampling during data generation hurts the downstream quality of programs synthesized by edit sequence models (Figures 5 and 6).
@@ 2 LintSeq: Code Synthesis as a Sequential Edit Problem
Figure 2: Visualizing LintSeq, an algorithm for re-factoring programs into sequences of edits. This algorithm samples across all of the static error-free line insertions that can be used to write a program chunk-by-chunk. It uses the Unix-diff operator to express generated edit sequences as text strings.
LintSeq is an algorithm for synthetic data generation that samples across sequences of insertions that can be used to sequentially write a program, leveraging a linter to repeatedly check program states for static errors during generation. A linter is a type of standard code analysis tool that verifies the correctness of programs. The LintSeq algorithm is loosely inspired by recent work on discrete diffusion methods for text generation, where decoding is non-autoregressive (Li et al., 2022a).
The key hypothesis underlying LintSeq is as follows: by training LMs on code edit sequences with teacher-forced supervised learning, we can potentially achieve a better trade-off between generation quality and compute at inference-time while still benefiting from the training and sampling efficiency of the autoregressive modeling paradigm. In this section, we provide a formalism for the edit sequence re-parameterization of code synthesis and we formally introduce LintSeq.
2.1 Reparameterizing Code Datasets with Edits
@@
Let $\Delta(\cdot,\cdot)$ denote the Unix diff operator (Thompson & Ritchie, 1975), which computes a text difference between a pair of strings by performing a line-by-line matching and returns a string summarizing the detected differences. The diff operator is implemented by popular version control and software development systems to help programmers track edits or code “diffs” between versions of text files. A single edit computed with the diff operator may consist of multiple line deletions and/or line insertions.
Fix a program $y$ in the dataset $\mathcal{D}$. Consider a sequence $\bm{\sigma}_{y}$ of $j$ text strings corresponding to programs or program states that terminates at $y$.
$\bm{\sigma}_{y}=(y_{1},\dots,y_{j-1},y)$ (1)
We can equivalently re-express $\bm{\sigma}_{y}$ as an edit sequence $\bm{\delta}_{y}$ of length $j$ by first computing a diff between an empty program $\epsilon$ and the first program in the sequence, and then computing diffs between all pairs of consecutive programs, as shown below.
@@
If $\mathcal{D}^{\prime}$ is a dataset such that for every pair $(x,y)\in\mathcal{D}$, there exists a pair $(x,\bm{\delta}_{y})\in\mathcal{D}^{\prime}$, then we say that $\mathcal{D}^{\prime}$ is an edit sequence re-factoring of $\mathcal{D}$.
Figure 3: Empirics of processing code data with LintSeq. Left: Lines per example in a dataset of instruction finetuning data for Python synthesis before and after processing with LintSeq via the linter pylint (see Section 3.2). LintSeq processing adds lines of diff metadata to examples (see Appendix A). Right: The corresponding edit counts per synthetic code edit sequence. On a dataset of short programs (14 lines of code, on average), the mean LintSeq edit sequence contains four edits.
@@ 2.2 Generating Linter-Guided Synthetic Edit Sequences
Recall from above that a single program edit computed by the diff operator $\Delta(\cdot,\cdot)$ can consist of any number of deletions and insertions. LintSeq is an algorithm for computing edit sequence re-factorings $\mathcal{D^{\prime}}$ such that all data $(x,\bm{\delta}_{y})\in\mathcal{D^{\prime}}$ have a particular property: every edit in $\bm{\delta}_{y}$ consists of insertions only. There are two phases in LintSeq: a backward sampling phase that is used to compute program state sequences $\bm{\sigma}_{y}$, and a forward edit sequence computation phase that is used to re-express $\bm{\sigma}_{y}$ as edit sequences $\bm{\delta}_{y}$. An illustration of these phases is shown in Figure 2. Full examples of edit sequences generated with LintSeq are provided in Appendix E (Figures 9 and 10).
@@
Once $s$ program state sequences $\bm{\sigma}_{y}$ have been generated for each $(x,y)\in\mathcal{D}$, we run the forward edit computation phase of our algorithm. In this phase, we apply Equation 2 from above to compute an edit sequence $\bm{\delta}_{y}$ for each $\bm{\sigma}_{y}$. Starting from the last program that was added to $\bm{\sigma}_{y}$, we use the diff operator to compute edits between each pair of consecutive programs in $\bm{\sigma}_{y}$ up to the original program $y$. Finally, we pair each edit sequence $\bm{\delta}_{y}$ with its instruction $x$ (if present) to yield an edit sequence re-factoring $\mathcal{D}^{\prime}$ of $\mathcal{D}$ with size $sN$.
@@ 2.3 Properties of LintSeq Data
Synthetic edit sequences generated by LintSeq have a few other important properties. Let $\bm{\delta}_{y}$ be an arbitrary $j$-length edit sequence in $\mathcal{D}^{\prime}$ generated with LintSeq, $\bm{\delta}_{y}=(\Delta(\varepsilon,y_{1}),\dots,\Delta(y_{j-1},y))$. First, we observe that there is a simple correspondence between $\bm{\delta}_{y}$ and the original program $y$ used to generate it: $y$ can be re-constructed by starting with an empty program, and successively applying each edit in $\bm{\delta}_{y}$ to this program one-by-one. In other words, the edit sequence $\bm{\delta}_{y}$ resolves to $y$. Furthermore, by construction, every prefix subsequence of $\bm{\delta}_{y}$ resolves to a sub-program of $y$ that is error-free, i.e. that throws no errors when checked with the linter or the verifier used during generation. These two properties, in conjunction with the uniform sampling step used in the first phase of the algorithm, show that LintSeq samples $s$ examples across all possible static error-free sequences of line insertions that can be used to sequentially write a program $y$ from-scratch.
We show an example of program synthesis dataset statistics before and after LintSeq processing in Figure 3. In the worst case, re-expressing a program as an edit sequence increases the length of a training example by a token count that is constant in the number of program lines (Appendix A).
@@ 2.4 Practicalities of Training Language Models on LintSeq Data
LintSeq can be run on any code data. It is agnostic to the contents of a program, and only depends on knowledge of the programming language that a file is written in and the existence of a linter or another kind of verifier for written program files.
We use teacher-forced supervised learning (Williams & Zipser, 1989) to train models on LintSeq data, concatenating edit sequences into a single string by interleaving edits with special tokens, “<|diff|>,” and computing instruction-conditioned losses over the resultant sequences. During inference, finetuned models can be prompted to synthesize programs with edit sequences by appending these special tokens to the ends of prompts. More details are provided in Appendix A.
Synthetic data generation with LintSeq is controlled by a single hyperparameter: the number of edit sequences $s$ that are sampled for each example in the source code dataset $\mathcal{D}$. Edit sequence sampling can optionally be constrained to avoid repetitions, though this may yield a dataset $\mathcal{D}^{\prime}$ with fewer than $s|\mathcal{D}|$ examples if there are programs in $\mathcal{D}$ with no non-empty and static error-free sub-programs.
@@ 3 Experiments
To study LintSeq and the impact of re-parameterizing program synthesis as a sequential edit generation problem, we conduct a multi-pronged set of instruction finetuning experiments. These experiments study code synthesis in Python and are designed to answer the following questions:
  • How does finetuning tiny LMs on re-factorized code edit data generated with LintSeq impact benchmark coverage compared to finetuning on the original code data?
  • Do performance improvements hold across model scales, families, and tokenizers?
  • How does ablating linter-guidance from LintSeq to finetune on randomly sampled edit sequences impact code synthesis?
Similar to previous works (Chen et al., 2021), we evaluate models by computing zero-shot coverage statistics on code synthesis benchmarks with and without repeated sampling. We expound upon our motivation for evaluating models in this manner in Appendix B.3.1.

### v3

@@ Abstract
Software engineers mainly write code by editing existing programs. In contrast, language models (LMs) autoregressively synthesize programs in a single pass. One explanation for this is the scarcity of sequential edit data. While high-quality instruction data for code synthesis is scarce, edit data for synthesis is even scarcer. To fill this gap, we develop a synthetic data generation algorithm called LintSeq. This algorithm refactors programs into sequences of synthetic edits by using a linter to procedurally sample across interdependent lines of source code. Synthetic edits sampled with LintSeq reflect the syntax and semantics of their programming language. To test the algorithm, we use it to refactor a dataset of instruction + program pairs into instruction + program-diff-sequence tuples. Then, we fine-tune a series of smaller LMs ranging from 2.6B to 14B parameters on both the re-factored and original versions of this dataset. We perform comprehensive evaluations comparing edit sequence code LMs against baselines on HumanEval, MBPP(+), CodeContests, DS-1000, and BigCodeBench. We show that models fine-tuned to iteratively synthesize code match or outperform baselines on pass@1, and exhibit better scaling across higher pass@k as a function of total test-time FLOPs. Finally, we also pretrain our own tiny LMs for code understanding. We show that fine-tuning these models to synthesize code edit-by-edit results in strong performance on HumanEval and MBPP(+) compared to existing code language models of similar scale such as CodeT5+, AlphaCode, and Codex.
@@ 1 Introduction
The successes of language models (LMs) are difficult to overstate. However, consistent and correct zero-shot generation in code synthesis remains out-of-reach for all but the largest models (Abdin et al., 2024; Groeneveld et al., 2024; Dubey et al., 2024). Compared to other reasoning tasks, this setting has two challenging properties, namely solutions are both long and structured.
@@
Figure 1: Code synthesis with LMs trained on synthetic code edit sequences. Left: An example generation from an LM trained to synthesize code as a stream of linter-error-free edits. Right: Training LMs to write code edit-by-edit by preprocessing instruction data for SFT with LintSeq improves test-time scaling laws during repeated sampling, i.e. the percentage of benchmark problems solved by any attempt (pass@k) as a function of total test-time FLOPs compared to training on standard data (see Appendix A.4). Shading indicates standard error in linear fit.
@@
In contrast, LMs are trained to autoregressively synthesize entire programs from scratch. This makes repeatedly editing a program with an LM extremely expensive – current state-of-the-art, LM-powered code editing tools like Cursor repeatedly prompt models to rewrite entire programs during every edit generation call (Sanger, 2024). LM outputs also suffer from degrading quality as sequence lengths grow and exhibit limited diversity across samples (Chen et al., 2021; Li et al., 2022b; Roziere et al., 2023; Lozhkov et al., 2024). The consequence of these pathologies is that there does not exist a reliable trade-off between zero-shot generation quality and total test-time compute under the current paradigm of autoregressive code synthesis, particularly for smaller LMs.
@@
In this paper, we claim that these issues can be mitigated at the data-level by reparameterizing code synthesis as a sequential edit problem. Rather than training models for single-step generation of entire programs, we propose that models be trained to generate code edit-by-edit. This objective has a major obstacle: while datasets of filtered GitHub repository commits like CommitPackFT (Muennighoff et al., 2023) have dramatically improved the quality of open-source code edit data, they contain limited sequential data. Moreover, the edits in such these datasets reflect the granularity at which programmers save code, but not necessarily the granularity at which they write and/or reason about it.
@@
To address this, we introduce a sampling algorithm called “LintSeq” that can be used to express any program in a training corpus as a sequence of structured code edits. LintSeq leverages linters – simple code analysis tools that check programs for errors and stylistic issues – to ensure that each generated edit meaningfully reflects the syntactical structure of the programming language that it is written in. The algorithm consists of two phases: a backward phase, which takes a source file as input and samples code deletions from this file to yield possible sequences of linter-error-free intermediate program states; and a forward edit computation phase, which reverses each sampled program sequence, employs the Unix diff (Thompson & Ritchie, 1975) operator to compute deltas between consecutive versions of each file, and outputs the generated edit sequences. LMs trained on data sampled with LintSeq synthesize code by repeatedly predicting insertion edits to files.
To test the impact of training LMs on synthetic edit sequences sampled with LintSeq, we conduct a series of supervised fine-tuning (SFT) experiments. In each experiment, we compare the performance of models trained on a corpus of example programs re-sampled into synthetic edit sequences with LintSeq to those trained on the original dataset. We evaluate LMs zero-shot and without chain-of-thought on HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), DS-1000 (Lai et al., 2023), BigCodeBench (Zhuo et al., 2024), and CodeContests (Li et al., 2022b) on “pass@k,” the proportion of problems solved by any attempt given “k” tries. Our results show the following:
@@
    Across models ranging in scale from 150M to 14B parameters, training LMs to iteratively synthesize programs improves the diversity of model-generated code compared to training on standard instruction data, while either preserving or improving code quality.
@@
    The improved diversity of generated programs means that pass@k performance increases faster as a function of test-time FLOPs, allowing for a better trade-off between the two.
@@
    Ablating the linter from edit sampling during data generation hurts the downstream quality of programs synthesized by edit sequence models.
@@ 2 LintSeq: Code Synthesis as a Sequential Edit Problem
Figure 2: LintSeq: Training LMs to write code edit-by-edit with supervised learning by generating synthetic data. LintSeq decomposes existing programs into synthetic edits that reflect the syntax & semantics of their programming language. At each iteration, the algorithm samples an edit chunk from a program by: randomly selecting a line of code to delete; identifying the minimal set of lines that are dependent on this line with a code linter; and finally, removing the line and its dependents. These steps are repeated until all lines of code have been removed. LintSeq then processes the reversed sequence of program states with Unix-diff to express it as a sequence of edits.
The key to solving a hard problem often lies in knowing how to decompose it into sub-problems. LintSeq is an algorithm for synthetic data generation that decomposes programs in training corpuses across insertion edit chunks that reflect the syntax and semantics of their programming language. To sample such chunks, it uses a code linter. The algorithm is inspired by recent work on discrete diffusion methods for text generation, where decoding is non-autoregressive (Li et al., 2022a).
Informally, the hypothesis underlying LintSeq is as follows: by training LMs to synthesize code edit-by-edit on large-scale datasets, we can potentially achieve a better trade-off between generation quality and test-time compute while still benefiting from the training and sampling efficiency of autoregressive language modeling. In this section, we define important terms, provide a formalism for the edit sequence re-parameterization of code synthesis, and formally introduce LintSeq.
2.1 Definitions
We define a linter to be a static code analysis tool that scans source code for defects. Linters can identify code that is objectively incorrect, throwing errors if and when a program contains syntax errors or refers to non-existent variables or packages. It is important to note that unlike a formal verifier, linters may return false positives, i.e. they may be unable to detect more complex errors, particularly in dynamically typed programming languages like Python or JavaScript.
For a given source file, define an intermediate program state to be a program that contains only a subset of the line-by-line contents of the original file, such that the order of these lines is preserved. We call an intermediate program state linter-error-free if checking this program with an appropriate linter produces exactly the same error trace(s) as those output when checking the original source file.
2.2 Representing Code with Edit Sequences
@@
Let $\Delta(\cdot,\cdot)$ denote the Unix diff operator (Thompson & Ritchie, 1975), which computes a text difference between a pair of strings by performing a line-by-line matching and returns a summary of the detected differences. The diff operator is implemented by popular version control and development systems to help programmers track edits between versions of text files. A single edit computed with the diff operator may consist of multiple line deletions and/or line insertions.
Fix a program $y$ in the dataset $\mathcal{D}$. Consider a sequence of $\bm{\sigma}_{y}$ of $j$ text strings corresponding to programs that terminates at $y$, $\bm{\sigma}_{y}=(y_{1},\dots,y_{j-1},y)$. We can equivalently re-express $\bm{\sigma}_{y}$ as an edit sequence $\bm{\delta}_{y}$ of length $j$ by first computing a diff between an empty program $\varepsilon$ and the first program in the sequence, and then computing diffs between all pairs of consecutive programs, as shown below.
@@
If $\mathcal{D}^{\prime}$ is a dataset such that for every pair $(x,y)\in\mathcal{D}$, there exists a pair $(x,\bm{\delta}_{y})\in\mathcal{D}^{\prime}$, then we say that $\mathcal{D}^{\prime}$ is an edit sequence refactoring of $\mathcal{D}$.
@@ 2.3 Generating Linter-Guided Synthetic Edit Sequences
Recall from above that a single program edit computed by the diff operator $\Delta(\cdot,\cdot)$ can consist of any number of deletions and insertions. LintSeq is an algorithm for computing edit sequence refactorings $\mathcal{D^{\prime}}$ such that all data $(x,\bm{\delta}_{y})\in\mathcal{D^{\prime}}$ have a particular property: every edit in $\bm{\delta}_{y}$ consists of insertions only. There are two phases in LintSeq: a backward sampling phase that is used to compute program state sequences $\bm{\sigma}_{y}$, and a forward edit sequence computation phase that is used to re-express $\bm{\sigma}_{y}$ as edit sequences $\bm{\delta}_{y}$. Pseudo-code as well as a visualization of each of these phases is provided in Figure 2. Full examples of edit sequences generated with LintSeq are provided in Appendix F (Figures 9 and 10).
@@
Once $s$ program state sequences $\bm{\sigma}_{y}$ have been generated for each $(x,y)\in\mathcal{D}$, we run the forward edit computation phase of our algorithm. In this phase, we apply Equation 1 from above to compute an edit sequence $\bm{\delta}_{y}$ for each $\bm{\sigma}_{y}$. Starting from the last program that was added to $\bm{\sigma}_{y}$, we use the diff operator to compute edits between each pair of consecutive programs in $\bm{\sigma}_{y}$ up to the original program $y$. Finally, we pair each edit sequence $\bm{\delta}_{y}$ with its instruction $x$ (if present) to yield an edit sequence refactoring $\mathcal{D}^{\prime}$ of $\mathcal{D}$ with size $sN$.
@@ 2.4 Properties of LintSeq Data
Synthetic edit sequences generated by LintSeq have a few other important properties. Let $\bm{\delta}_{y}$ be an arbitrary $j$-length edit sequence in $\mathcal{D}^{\prime}$ generated with LintSeq, $\bm{\delta}_{y}=(\Delta(\varepsilon,y_{1}),\dots,\Delta(y_{j-1},y))$. First, we observe that there is a simple correspondence between $\bm{\delta}_{y}$ and the original program $y$ used to generate it: $y$ can be re-constructed by starting with an empty program, and successively applying each edit in $\bm{\delta}_{y}$ to this program one-by-one. In other words, the edit sequence $\bm{\delta}_{y}$ resolves to $y$. Furthermore, by construction, every prefix subsequence of $\bm{\delta}_{y}$ resolves to a intermediate program state of $y$ that is linter-error-free (see Section 2.1). These two properties, in conjunction with the uniform sampling step used in the first phase of the algorithm, show that LintSeq samples $s$ examples across all possible linter-error-free sequences of line insertions that can be used to sequentially write a program $y$ from-scratch.
We show an example of program synthesis dataset statistics before and after LintSeq processing in Appendix A (Figure 6). In the worst case, re-expressing a program as an edit sequence increases the length of a training example by a token count that is constant in the number of program lines [^1] See Appendix B for more details. .
@@ 2.5 Practicalities of Training Language Models on LintSeq Data
LintSeq can be run on any code data. It is agnostic to the contents of a program, and only depends on knowledge of the language that a program is written in, and the existence of a linter for this language.
We use teacher-forced supervised learning (Williams & Zipser, 1989) to train models on LintSeq data, concatenating edit sequences into a single string by interleaving edits with special tokens, “<|diff|>,” and computing instruction-conditioned losses over the resultant sequences. At test-time, fine-tuned models can be prompted to synthesize programs with edit sequences by appending these special tokens to the ends of prompts. More details are provided in Appendix B.
Synthetic data generation with LintSeq is controlled by a single hyperparameter: the number of edit sequences $s$ that are sampled for each example in the source code dataset $\mathcal{D}$. Edit sequence sampling can be constrained to avoid repetitions.
@@ 3 Experiments
To study LintSeq and the impact of re-parameterizing program synthesis as a sequential edit generation problem, we conduct a set of supervised fine-tuning (SFT) experiments. These experiments study code synthesis in Python and are designed to answer the following questions:
  • How does fine-tuning tiny code LMs to generate programs edit-by-edit with supervised learning impact performance on benchmarks compared to fine-tuning on standard code data?
  • Do performance improvements hold for “off-the-shelf” LMs and on harder coding benchmarks? Do they hold across model scales, tokenizers, and families?
  • How does ablating linter-guidance from LintSeq impact test-time performance?
Similar to previous works (Chen et al., 2021), we evaluate models by computing “pass@k,” the probability that at least one of “k” generations for a problem passes all of the unit tests.